# Chain of Summaries: Summarization Through Iterative Questioning

### v1

@@ 5 Results
Table 2: Improvement in F1-score of Chain-of-Summaries against the zero-shot summarization baseline on TriviaQA. Memorization lists the performance on those questions based on which the summary was created (either from the human-crafted set or based on synthetic questions). Generalization lists the performance on the validation set of TriviaQA. Synthetic questions resemble a viable alternative when no ground-truth questions are available.
@@
We investigated three key hyperparameters in our CoS approach: number of iterations, questions per iteration, and integration of Chain of Draft (CoD) a prompting method where model is using reduced number of thinking output tokens. As shown in Appendix C, distributing questions over multiple iterations consistently outperforms presenting all questions simultaneously. For Llama3.2:3B, using 10 iterations with 1 question per iteration improved generalization from 0.52 to 0.64, while the single iteration approach with 50 questions reduced performance to 0.49. Similarly, for Qwen2.5:7B, the 10-iteration configuration improved generalization from 0.48 to 0.58, while the single-iteration approach reduced performance to 0.39. The inclusion of CoD showed model-dependent effects. For Llama3.2:3B, CoD achieved the highest generalization score of 0.68 with 10 iterations and 1 question per iteration - an improvement of 6%. However, for Qwen2.5:7B, no CoD performed better (0.60 versus 0.56). These results highlight the importance of balancing the number of iterations with the number of questions per iteration, and taking into account model-specific characteristics.
6 Fine-tuning Analyses
To further understand whether the iterative multi-step process can be compressed into a single step with an efficient language model, we fine-tune a reward model to estimate downstream QA performance, and then use GRPO to fine-tune a small Qwen model for summarization.
Fine-tuning ModernBERT for Downstream QA Performance Estimation
For the reward model, we fine-tuned a AnswerDotAI/ModernBERT-base model warner2024smarterbetterfasterlonger with a context size of 8,192 using CoS intermediate and final summaries paired with their respective attained F1 scores. A detailed description of the training dataset is available in Appendix D. We employed binary cross-entropy loss with sigmoid activation to match the predicted distribution directly with the targets (range: 0–1). Training the model with a batch size of 16, gradient accumulation steps of eight, and a learning rate of 0.008 over 15 epochs produced a model with training and validation losses of 0.434 and 0.508, respectively, and a test RMSE of 0.147, which demonstrates its strong predictive capability. This enables the Chain of Summaries process to be completed more quickly and allows existing summaries to be evaluated and selected for refinement if necessary.
Fine-tuning Qwen for summarization
We explored reducing our multi-step CoS methodology into a single-step operation by reinforcement fine-tuning Qwen3:0.6B and Qwen2:0.5B-Instruct using GRPO guo2025deepseek; shao2024deepseekmathpushinglimitsmathematical with our ModernBERT estimator as reward model. The reward function was formulated as |1s^|-|1-$\hat${s}|, where $\hat{s}$ represents the predicted downstream QA performance score, creating a gradient toward information-dense representations. We used 4 samples per batch with 4 candidate generations trained for three epochs with BF16 precision. The prompt template is provided in Appendix A. The dataset of (interediate) summaries and their attained scores is provided in Appendix D.
Despite their small size, our fine-tuned smaller Qwen models produced strong results as shown in Table 1. Fine-tuned Qwen2:0.5B-Instruct achieved a TriviaQA generalization score of 0.74 (GPT-4o-mini backbone), approaching the 0.80 of the larger CoS approach. In the Llama3.2:3B evaluation, the fine-tuned Qwen3:0.6B model matched the full-source performance on TriviaQA (0.50 vs. 0.51) and SQUAD1.1 (0.19 vs. 0.19). While these results demonstrate that small models can produce high-quality summaries that are comparable to those of larger approaches, the significant discrepancy on TruthfulQA (5.33 vs. 14.77) suggests that the dialectical refinement of CoS captures nuances that are difficult to distill into a single, compact model.
Downstream QA task comparison
@@ 6 Fine-tuning Analyses
Table 3: TriviaQA F1 scores. Our fine-tuning-free CoS method with a GPT-4o-mini backbone achieves 80.0% F1, outperforming strong retrieval-augmented baselines like ATLAS (74.5%) and REPLUG (77.3%).
7 Discussion
@@
Our main finding is that CoS creates a textual summary that is more useful to a downstream LLM than the source text by condensing information and removing irrelevant context – suggesting that CoS makes information better accessible for downstream language model use. Moreover, CoS consistently outperforms zero-shot summarization, as well as dedicated summarization methods such as BRIO and PEGASUS.
@@
Our findings are in agreement with recent work showing that increased test-time compute (here, multiple LLM calls) can improve the quality of generated text muennighoff2025s1 – which would amortize over multiple requests in our envisioned scenario, where website maintainers would use CoS to create information-dense summaries for downstream LLM agents to process. If implemented server-side, which is our primary recommendation, website maintainers can inspect the current summary of their website and/or specific subpages, and ingest manual corrections if needed. If they don’t have human-crafted Q&A pairs, website maintainers can use synthetic Q&A pairs, which we have shown to perform on par with real Q&A pairs. In contrast, client-side implementation (e.g., on the site of an LLM agent operating on the web) would induce redundancy as different systems would have to maintain their caches – and website maintainers would have little control over how their websites are summarized and whether the cache may have been invalidated. Our attempt at fine-tuning an LLM towards good one-step summarization was less effective than the multi-step summarization. We cannot exclude that this would be different with a larger model or through a more sophisticated reward model, which would change this result. However, as of now, we can conclude from this that the iterative procedure is necessary to obtain a high-quality summary. In sum, we have shown how Hegel’s dialectical method can be leveraged with large language models to obtain a general-purpose summary that anticipates future questions, ready to be used as an LLM-friendly cache for content on the web.
8 Conclusion
@@
9 Limitations
Our selection of datasets mainly consists of question-answering datasets. While we deem this a reasonable proxy for practical applications, it does not involve complex multi-stage synthesis of content on the web from various sources, which would need to be carried out by the client LLM. However, summarizing multiple possibly interlinked documents on the web for LLM use would be an exciting avenue for future work, testing how such links within the summaries could facilitate navigation of LLMs on the web. Future work may further look at more sophisticated question stratification and fine-tuning larger LLMs.
10 Ethical Considerations
We don’t expect that the proposed method would cause any extra harm compared to LLMs operating on the web on their own. In contrast, if implemented server-side as an LLMs.txt file, our approach allows website maintainers to validate what textual representation LLMs process.
@@ Appendix A Prompts
"content": "Given a question and content, answer the question as simply as possible! Dont answer in complete sentences; words and phrases are sufficient. Answer the question based on the content provided. If the answer is not present in the content, say \"I dont know.",
@@
k}
Listing 4: Prompt template for GRPO finetuning
template = """
TITLE: {title}
CONTENT: {content}
SUMMARY:"""
Appendix B Summary Length vs. Generalization
Our analysis reveals a consistent correlation between summary length and model performance. As shown in Table 4, Chain of Summaries achieves the optimal balance between information density and token efficiency across all model architectures. The token efficiency advantage is most pronounced when comparing CoS with content-based approaches, which require approximately 65x more tokens (11,219.68 vs. 170.75 for GPT-4o-mini) while achieving lower F1 scores. This pattern holds for models of all sizes, from Qwen2.5 to GPT-4o-mini, indicating that the benefits of CoS are not merely an artifact of model size. Regarding computational efficiency for downstream applications, CoS summaries require substantially fewer tokens while preserving critical information. This allows for more efficient utilization of limited context windows in retrieval-augmented systems.
Model Metric Brio Content CoS Finetune-GRPO Pegasus Zero-Shot
GPT-4o-mini F1 Score 0.6303 0.7569 0.7981 0.7296 0.7214 0.7307
Token Count 117.00 11219.68 170.75 330.68 125.62 76.63
Llama3.2:3B F1 Score 0.4460 0.5053 0.6499 0.4939 0.4965 0.4947
Token Count 117.00 11219.68 223.19 330.68 125.62 251.00
Qwen2.5 F1 Score 0.2067 0.4057 0.5719 0.4322 0.4298 0.4787
Token Count 117.00 11219.68 90.40 330.68 125.62 75.95
Table 4: Performance Comparison of summarization Methods on TriviaQA Dataset. The table presents F1 scores and token counts across six summarization approaches with three model backbones, where token counts and F1 scores are the mean across multiple runs. Chain of Summaries (CoS) consistently outperforms other methods, achieving higher F1 scores with reasonable token efficiency. Bold values indicate best performance per model. While content-based approaches maintain competitive F1 scores, they require significantly more tokens, highlighting CoS’s ability to create information-dense summaries that effectively support question answering tasks.
Appendix C Hyperparameter Sensitivity Analysis
Our Chain of Summaries approach can be configured with different hyperparameters that impact performance. Table 5 provides a detailed analysis of the effect of varying the number of iterations and questions per iteration on the performance of the Llama3.2:3B and Qwen2.5:7B models. Several key findings emerge from this analysis: Distributing questions across multiple iterations outperforms a single-batch presentation for both the tested models. Llama3.2:3B achieves optimal generalization (0.68) with Chain of Drafts and 10 iterations/question, whereas Qwen2.5:7B achieves its best performance (0.60) without Chain of Drafts using the same configuration. The five iterations/10 questions approach maximizes memorization scores, but more gradual refinement (10 iterations/one question) better supports generalization. The effectiveness of CoD varies by architecture, benefiting Llama3.2:3B (+6.3%) while Qwen2.5:7B performs optimally without it. The results emphasize the importance of model-specific hyperparameter tuning for Chain of Summaries implementations.
Method T Metrics S Metrics Configuration
Llama3.2:3B
Chain of Summaries w/ CoD 0.21 $\rightarrow$ 0.29 0.48 $\rightarrow$ 0.45 0.21 $\rightarrow$ 0.30 0.48 $\rightarrow$ 0.44 1 50
Chain of Summaries w/ CoD 0.21 $\rightarrow$ 0.36 0.48 $\rightarrow$ 0.61 0.20 $\rightarrow$ 0.34 0.48 $\rightarrow$ 0.62 5 10
Chain of Summaries w/ CoD 0.21 $\rightarrow$ 0.33 0.48 $\rightarrow$ 0.68 0.20 $\rightarrow$ 0.33 0.48 $\rightarrow$ 0.67 10 1
Qwen2.5:7B
Chain of Summaries w/ CoD 0.23 $\rightarrow$ 0.49 0.47 $\rightarrow$ 0.36 0.23 $\rightarrow$ 0.51 0.46 $\rightarrow$ 0.37 1 50
Chain of Summaries w/ CoD 0.23 $\rightarrow$ 0.59 0.47 $\rightarrow$ 0.47 0.23 $\rightarrow$ 0.58 0.46 $\rightarrow$ 0.47 5 10
Chain of Summaries w/ CoD 0.23 $\rightarrow$ 0.41 0.47 $\rightarrow$ 0.54 0.23 $\rightarrow$ 0.41 0.46 $\rightarrow$ 0.56 10 1
Table 5: Hyperparameter Sensitivity Analysis of Chain of Summaries Approach on TriviaQA Dataset. This table demonstrates how different configurations of iterations and questions per iteration (Q per I) affect model performance across Llama3.2:3B and Qwen2.5:7B models. Performance is measured using training metrics (T Metrics) and validation metrics (S Metrics), each reporting memorization and generalization capabilities. The results reveal that distributing questions across multiple iterations (10 iterations with 1 question or five iterations with 10 questions) consistently outperforms concentrating all questions in a single iteration (1 iteration with 50 questions). For Llama3.2:3B, Chain of Summaries with Chain of Draft (CoD) achieves the highest generalization scores (0.68/0.67) when using 10 iterations with 1 question, while Qwen2.5:7B performs best without CoD. Bold values indicate best overall performance, and italic values show second-best performance for each model and metric combination.
Appendix D Data Appendix
This appendix provides details of the dataset created to train the reward model for our Chain of Summaries (CoS) framework. The dataset comprises machine-generated summaries, each paired with an F1 score that quantifies its usefulness in subsequent question-answering tasks. This data was synthetically generated as an output of our iterative CoS procedure, in which candidate summaries were produced by applying our model to a diverse corpus of documents. These documents originate from TriviaQA, TruthfulQA, and SQUADv1.1. Each summary was then automatically evaluated against a set of validation questions from its source document that was provided in the dataset, and an F1 score was computed based on the proportion of questions that could be answered correctly using only the information within the summary. This process yielded a collection of (summary, F1_score) pairs. The dataset is structured for a regression task where the model learns to predict the F1 score, a continuous value in the range $[0,1]$, from the summary text. The resulting reward model can thereby estimate a summary’s capabilities for the QA task.
Preprocessing
To ensure the quality and suitability of the dataset, we applied a number of pre-processing steps. We filtered out any malformed entries or entries with missing F1 scores, and we performed deduplication based on the summary text. This meant that we retained only a single instance of any identical summaries. To guarantee that every entry could be processed by the reward model (AnswerDotAI/ModernBERT-base) without truncation, we removed all summaries that exceeded its context window of 8,192 tokens.
D.1 Characteristics
The training dataset comprises 17,971 entries. The key descriptive statistics for this dataset are presented in the Table below.
Statistic Value
Overall Dataset
Number of Entries 17,971
Summary Text Statistics
Mean Length (tokens) 232.60
Median Length (tokens) 146.0
Max Length (tokens) 23756
Min Length (tokens) 6
F1 Score Statistics
Mean 0.404
Median 0.333
Standard Deviation 0.308
Range [0.0, 1.0]
Table 6: Descriptive statistics of the reward model training dataset.
Figure 3 shows how summary lengths are distributed across the dataset. There is a clear pattern where shorter summaries are the most frequent, with a large proportion of the data concentrated below the 1,000-token threshold. As summaries become longer, they occur less frequently. This distribution ensures that, while the dataset contains many concise examples, it also captures a broad range of textual complexity by including longer, more detailed summaries.
Figure 3: Distribution of summary lengths, measured in gpt-4o-mini tokens.
As shown in Figure 4, the quality score distribution is well-suited for training a robust reward model. The large populations of summaries that either answer none of the questions (a score of 0.0) or answer all of them (a score of 1.0) provide clear negative and positive anchors. Furthermore, the distinct clusters at intermediate scores, representing summaries that answer only some of the required questions, are crucial for teaching the model to discern nuanced differences in utility rather than simply making a binary ’good’ or ’bad’ judgment.
Figure 4: Distribution of F1 scores in the training dataset used for the reward model.

### v2

@@ 5. Results
Table 3. Improvement in F1-score of Chain-of-Summaries against the zero-shot summarization baseline on TriviaQA. Memorization lists the performance on those questions based on which the summary was created (either from the human-crafted set or based on synthetic questions). Generalization lists the performance on the validation set of TriviaQA. Synthetic questions resemble a viable alternative when no ground-truth questions are available.
Table 4 contrasts representative synthetic questions produced by our generator with human-authored questions drawn from the TriviaQA evaluation set. The examples are from the same topical neighborhood (Sherlock Holmes) to highlight qualitative differences attributable to question provenance rather than subject matter.
Synthetic (Generated) Human (TriviaQA Eval)
Who created the character of Sherlock Holmes?
Answer: Arthur Conan Doyle
In the Sherlock Holmes stories who was Moriarty’s second in command?
Answer: Colonel Moran
In what year did Sherlock Holmes first appear in print?
Answer: 1887
What item of headwear is associated with Sherlock Holmes?
Answer: Deerstalker hat
Where is Holmes’s address located?
Answer: 221B Baker Street
In which London magazine did Sherlock Holmes first appear?
Answer: The Strand
Table 4. Side-by-side examples of synthetic versus human questions in the Sherlock Holmes domain. Synthetic questions tend to be more canonical and factoid-centric, whereas human questions more often reflect culturally mediated associations and narrative-specific details.
@@
We investigated three key hyperparameters in our CoS approach: number of iterations, questions per iteration, and integration of Chain of Draft, a prompting method where the model uses a reduced number of thinking output tokens. As shown in Table 5, distributing questions over multiple iterations consistently outperforms presenting all questions simultaneously. For Llama3.2:3B, using 10 iterations with 1 question per iteration improved generalization from 0.52 to 0.64, while the single iteration approach with 50 questions reduced performance to 0.49. Similarly, for Qwen2.5:7B, the 10-iteration configuration improved generalization from 0.48 to 0.58, while the single-iteration approach reduced performance to 0.39. The inclusion of Chain of Draft showed model-dependent effects. For Llama3.2:3B, Chain of Draft achieved the highest generalization score of 0.68 with 10 iterations and 1 question per iteration - an improvement of 6%. However, for Qwen2.5:7B, no Chain of Draft performed better (0.60 versus 0.56). These results highlight the importance of balancing the number of iterations with the number of questions per iteration, and taking into account model-specific characteristics.
Method T Metrics S Metrics Configuration
Llama3.2:3B
Chain of Summaries w/ Chain of Draft 0.21 $\rightarrow$ 0.29 0.48 $\rightarrow$ 0.45 0.21 $\rightarrow$ 0.30 0.48 $\rightarrow$ 0.44 1 50
Chain of Summaries w/ Chain of Draft 0.21 $\rightarrow$ 0.36 0.48 $\rightarrow$ 0.61 0.20 $\rightarrow$ 0.34 0.48 $\rightarrow$ 0.62 5 10
Chain of Summaries w/ Chain of Draft 0.21 $\rightarrow$ 0.33 0.48 $\rightarrow$ 0.68 0.20 $\rightarrow$ 0.33 0.48 $\rightarrow$ 0.67 10 1
Qwen2.5:7B
Chain of Summaries w/ Chain of Draft 0.23 $\rightarrow$ 0.49 0.47 $\rightarrow$ 0.36 0.23 $\rightarrow$ 0.51 0.46 $\rightarrow$ 0.37 1 50
Chain of Summaries w/ Chain of Draft 0.23 $\rightarrow$ 0.59 0.47 $\rightarrow$ 0.47 0.23 $\rightarrow$ 0.58 0.46 $\rightarrow$ 0.47 5 10
Chain of Summaries w/ Chain of Draft 0.23 $\rightarrow$ 0.41 0.47 $\rightarrow$ 0.54 0.23 $\rightarrow$ 0.41 0.46 $\rightarrow$ 0.56 10 1
Table 5. Hyperparameter Sensitivity Analysis of Chain of Summaries Approach on TriviaQA Dataset. This table demonstrates how different configurations of iterations and questions per iteration (Q per I) affect model performance across Llama3.2:3B and Qwen2.5:7B models. Performance is measured using training metrics (T Metrics) and validation metrics (S Metrics), each reporting memorization and generalization capabilities. The results reveal that distributing questions across multiple iterations (10 iterations with 1 question or five iterations with 10 questions) consistently outperforms concentrating all questions in a single iteration (1 iteration with 50 questions). For Llama3.2:3B, Chain of Summaries with Chain of Draft achieves the highest generalization scores (0.68/0.67) when using 10 iterations with 1 question, while Qwen2.5:7B performs best without Chain of Draft. Bold values indicate best overall performance, and italic values show second-best performance for each model and metric combination.
Downstream QA Performance on TriviaQA
@@
Table 6. TriviaQA F1 scores. Our fine-tuning-free CoS method with a GPT-4o-mini backbone achieves 80.0% F1, outperforming strong retrieval-augmented baselines like ATLAS (74.5%) and REPLUG (77.3%).
6. Efficiency & Cost Analysis
CoS consistently outperforms traditional methods while using substantially fewer tokens (Figure 2). With only 170 tokens, GPT-4o-mini achieves F1=0.80, outperforming zero-shot (0.73), BRIO (0.63), and PEGASUS (0.72). Content-based approaches require over 11,000 tokens for F1=0.76, making them impractical. Llama3.2:3B reaches F1=0.62, exceeding zero-shot (0.52) with similar token efficiency. Qwen2.5 demonstrates the strongest efficiency gains, achieving F1=0.60 using just 42–47 tokens.
However, these efficiency gains come with an important drawback: the CoS approach requires an initial pre-processing step to create the summary structure. Generating a CoS summary incurs a one-time indexing cost that varies with document size. Table 7 presents a detailed breakdown of token costs across three representative document sizes, split by input and output tokens.
Phase Tokens 5k Doc 10k Doc 50k Doc
A: Question Generation Input 5,200 10,200 50,200
Output 300 300 300
B: Initial Summary Input 5,200 10,200 50,200
Output 100 170 500
C: Refinement (10 iter.) Input 57,150 108,550 515,150
Output 3,000 3,700 7,000
Total Indexing Input 67,550 128,950 615,550
Output 3,400 4,170 7,800
Summary Size 100 170 500
Savings per Query 4,900 9,830 49,500
Break-even (queries) $\sim$15 $\sim$14 $\sim$13
Indexing Cost (USD) \$0.012 \$0.022 \$0.097
Table 7. Token costs for CoS indexing across document sizes, split by input and output tokens. USD cost calculated using GPT-4o-mini pricing (\$0.15/M input, \$0.60/M output). Summary sizes scale proportionally with document length.
Consider a typical 10,000-token document. Phase A (Antithesis) consumes 10,200 input tokens (10,000 document + 200 prompt) and generates 300 output tokens (20 questions $\times$ 15 tokens each). Phase B (Thesis) requires 10,200 input tokens (document + prompt) and produces a 170-token summary. The cost is primarily dominated by Phase C (Synthesis) across 10 refinement iterations. Each iteration consists of two steps: (1) an answering step involving reading the current summary (170 tokens), the question list (300 tokens) and the prompt (200 tokens) to produce answers (200 tokens); and (2) a refining step involving reading the document (10,000 tokens), the current summary (170 tokens) and one unanswered question (15 tokens) to generate an updated summary (170 tokens). This yields approximately 10,855 input tokens and 370 output tokens per iteration, totalling 108,550 input tokens and 3,700 output tokens for 10 iterations.
The total indexing cost of approximately 133,000 tokens (128,950 input + 4,170 output) is offset by substantial inference savings: each subsequent retrieval using the 170-token CoS summary instead of the full 10,000-token document saves 9,830 tokens, a 98.3% reduction. The break-even point occurs after approximately 14 queries ($133{,}000\div 9{,}830\approx 13.5$). As Table 7 shows, this break-even ratio remains remarkably stable across document sizes (12-15 queries), making CoS “token positive” early in the deployment lifecycle for documents accessed repeatedly in retrieval-augmented systems.
Cost Implications
Using GPT-4o-mini API pricing (\$0.15 per million input tokens, \$0.60 per million output tokens), the per-query cost for a 10k document drops from \$0.0015 (full document) to \$0.000026 (CoS summary), a 98.3% reduction. For a document retrieved 1,000 times, CoS reduces inference costs from \$1.50 to \$0.026, saving \$1.45 after accounting for the one-time indexing cost of \$0.022. This efficiency gain scales multiplicatively across large document collections: indexing 1,000 documents costs approximately \$22, while saving over \$1,400 across 1,000 queries per document.
7. Discussion
@@
Our main finding is that CoS creates a textual summary that is more useful to a downstream LLM than the source text by condensing information and removing irrelevant context, suggesting that CoS makes information better accessible for downstream language model use. Moreover, CoS consistently outperforms zero-shot summarization, as well as dedicated summarization methods such as BRIO and PEGASUS.
@@
Our findings are in agreement with recent work showing that increased test-time compute (here, multiple LLM calls) can improve the quality of generated text (Muennighoff et al., 2025) which would amortize over multiple requests in our envisioned scenario, where website maintainers would use CoS to create information-dense summaries for downstream LLM agents to process. If implemented server-side, which is our primary recommendation, website maintainers can inspect the current summary of their website and/or specific subpages, and ingest manual corrections if needed. If they don’t have human-crafted Q&A pairs, website maintainers can use synthetic Q&A pairs, which we have shown to perform on par with real Q&A pairs. In contrast, client-side implementation (e.g., on the site of an LLM agent operating on the web) would induce redundancy as different systems would have to maintain their caches and website maintainers would have little control over how their websites are summarized and whether the cache may have been invalidated.
The computational demands of large language models carry environmental costs (Strubell et al., 2019). By reducing per-query token consumption by over 98%, CoS directly decreases inference energy requirements. For repeatedly-accessed documents, CoS’s amortized computational footprint is substantially lower than processing full documents. This reduction translates to lower carbon emissions (Schwartz et al., 2020), while enabling lower latency responses for real-time applications. CoS thus achieves greener AI by reducing inference compute while maintaining downstream task performance.
In sum, we have shown how Hegel’s dialectical method can be leveraged with large language models to obtain a general-purpose summary that anticipates future questions, ready to be used as an LLM-friendly cache for content on the web.
Limitations
Our selection of datasets mainly consists of question-answering datasets. While we deem this a reasonable proxy for practical applications, it does not involve complex multi-stage synthesis of content on the web from various sources, which would need to be carried out by the client LLM. Additionally, our current evaluation focuses exclusively on single-document summarization; future work should explore multi-document scenarios where iterative reasoning could consolidate knowledge across related documents. More broadly, summarizing multiple possibly interlinked documents on the web for LLM use would be an exciting avenue for future work, testing how such links within the summaries could facilitate navigation of LLMs on the web. Future work may further look at more sophisticated question stratification and fine-tuning larger LLMs.
8. Conclusion
@@
9. Ethical Considerations
We don’t expect that the proposed method would cause any extra harm compared to LLMs operating on the web on their own. In contrast, if implemented server-side, our approach allows website maintainers to validate what textual representation LLMs process.
@@ Appendix A Prompts
"content": "Given a question and content, answer the question as simply as possible! Dont answer in complete sentences; words and phrases are sufficient. Answer the question based on the content provided. If the answer is not present in the content, say \"I dont know.\"",